# SEP: Self-Enhanced Prompt Tuning for Visual-Language Model

### v1

@@ Abstract
Prompt tuning based on Context Optimization (CoOp) effectively adapts visual-language models (VLMs) to downstream tasks by inferring additional learnable prompt tokens. However, these tokens are less discriminative as they are independent of the pre-trained tokens and fail to capture input-specific knowledge, such as class-aware textual or instance-aware visual knowledge. Leveraging the discriminative and generalization capabilities inherent in pre-trained tokens, we introduce a novel approach named Self-Enhanced Prompt Tuning (SEP). The core principle of SEP involves adapting the learnable prompt tokens at each encoder layer from the corresponding self-pretrained tokens, thereby explicitly incorporating discriminative prior knowledge to enhance both textual-level and visual-level embeddings. Furthermore, SEP’s self-enhanced tokens not only boost discrimination but also mitigate domain shifts in unseen domains, enhancing generalization. In practice, SEP selects several representative tokens from all pre-trained tokens for each input data at every layer of the text/visual encoders. Subsequently, a Token Fusion Module (TFM) is introduced to generate a self-enhanced token by merging these representative tokens with the learnable tokens using a cross-attention mechanism. This self-enhanced token is then concatenated with all pre-trained tokens, serving as input for subsequent encoder layers to produce the relevant embeddings. Comprehensive evaluations across various benchmarks and tasks confirm SEP’s efficacy in prompt tuning. Code: https://github.com/htyao89/SEP.
@@ 1 Introduction
Recent studies demonstrate that the Visual Language Model (VLM) possesses a robust generalization capability in a variety of downstream tasks [2; 33]. However, the direct application of VLMs in scenarios with limited images is challenging because of their extensive parameters and the substantial image datasets required for training. Prompt tuning has recently emerged as an effective and straightforward method for leveraging the general knowledge of VLMs in specialized computer vision and machine learning tasks such as zero-shot learning, domain adaptation, domain generalization, and few-shot learning.
Prompt tuning combines additional learnable tokens with pre-trained ones to generate a discriminative embedding, termed Context Optimization (CoOp) [52]. Recent advances in CoOp-based methods [52; 45; 4; 31; 17; 5; 22; 18; 41; 54] aim to develop input-relevant prompt tokens to enhance the discriminative power of embeddings, as illustrated in Figure 1(a). The term ‘input-irrelevant prompt’ refers to a prompt initialized random that does not depend on specific input-related knowledge. However, since these prompts are derived from the training domain, they may exhibit domain shifts from unseen domains, potentially degrading performance. In addition, image-conditional prompts, which incorporate image features into the learnable tokens, have been proposed in [51; 48] (see Figure 1(b)). Despite their specificity, the image-conditional prompts have a limited ability to enhance the class-level classifier. In summary, while existing CoOp-based methods generate learnable tokens that are independent of input knowledge, they struggle to utilize essential input-aware prior knowledge, such as instance-aware visual knowledge or class-aware textual knowledge. Therefore, how to incorporate discriminative and generalizable input-aware knowledge into learnable prompts is crucial to enhancing their effectiveness for prompt tuning.
@@ 2 Related Works
Vision-Language Models: Recently, Vision-Language Models (VLMs) [33; 2], which utilize both visual and textual encoders, have been trained on a large-scale dataset of image-text pairs, demonstrating significant generalization and discrimination capabilities. To further enhance the descriptive power of VLMs, improvements have been made in several areas: (1) employing more robust text or visual encoders [42; 49; 25]; (2) integrating visual and textual knowledge more deeply [24; 39]; and (3) increasing the volume of images used [33; 16; 36; 37]. To enhance the diversity of text descriptions, Masked Language Modeling (MLM) [20; 27] has been employed, which randomly omits words in the text to aid in representation learning. In contrast to MLM, masked autoencoder-based methods [13] have been introduced to improve descriptive capabilities by randomly masking image patches. Among the VLM frameworks, CLIP stands out as a straightforward yet effective model that uses contrastive loss to train separate visual and text encoders based on 400 million image-text pairs. Given its robust generalization, many CoOp-based methods have been developed using CLIP to adapt pre-trained VLMs for downstream tasks. In this work, we have implemented a prompt-tuning strategy to generate task-specific embeddings.
@@
Prompt Tuning: To adapt pre-trained Vision-Language Models (VLMs) to downstream tasks, various strategies involving prompt tuning have been employed [10; 31; 22; 48; 33; 23]. Initially, CLIP[33] uses a hand-crafted template ‘a photo of a [CLASS]’ for zero-shot prediction. However, these hand-crafted prompts often fail to adequately capture the nuances of downstream tasks, leading to the development of textual prompt tuning strategies that use learnable textual tokens alongside class tokens to enhance textual embeddings. For example, the hand-crafted prompts are replaced with the learnable prompt tokens in Context Optimization(CoOp) [52]. Moreover, Conditional Context Optimization(CoCoOp) [51] and VPT [48] generate an image-conditional prompt fusing the image feature and the learnable textual prompt to boost generalization in the unseen domain. To better use the essential knowledge contained in frozen CLIP, Knowlege-Guided Context Optimization(KgCoOp) [45], ProGrad [53], and Prompt Regularization(ProReg) [54] all construct a novel constraint term by constraining the consistency between the learnable knowledge and the essential general knowledge. Unlike the above methods, which consider textual prompts, Ensembling Context Optimization(ECO) [1] employs prompt ensembling to combine multiple prompts. Furthermore, ProDA [28] considers the prompt’s prior distribution learning, and Distribution-Aware Prompt Tuning (DAPT) [5] optimizes the learnable prompt by maximizing inter-dispersion. Knowlege-Aware Prompt Tuning(KAPT) [17] employs external knowledge to boost generalization on unseen classes. PLOT [4] applies optimal transport to match the vision and text modalities for generating the discriminative and visual-aligned local textual prompt tokens. Besides the textual prompt tuning, Multi-modal Prompt Learning (MaPLe) [18] and PromptSRC [19] conduct the visual-textual prompt tuning by jointly conducting the prompt tuning on the visual and text encoders. Multitask Vision-Language Prompt Tuning (MVLPT) [38] integrates cross-task knowledge, enhancing the versatility of prompt tuning in VLMs. Finally, DenseCLIP [34] and CLIP-Adapter [12] introduce context-aware strategies and adapters to refine the embeddings for specific tasks, further illustrating the dynamic evolution of prompt tuning in the field of vision-language models. Recently, GraphAdaper [26] is proposed to fuse visual and textual knowledge with two Graph-based Adapters, and the Class-aware prompt is proposed in TCP [46] to inject class-level knowledge into the prompt.
@@ 3.1 Preliminaries
The frozen Contrastive Language-Image Pre-training(CLIP) trained with a large number of images exhibits robust generalization capacity to new classes or images, e.g., CLIP shows impressive performance on zero-shot learning and domain generalization tasks, Therefore, existing prompt tuning based on Context Optimization(CoOp) is proposed based on CLIP. Given an image along with all class names, CLIP extracts the visual and text embeddings with the visual and text encoders, and the contrast loss is calculated to align those two embeddings. To effectively adapt CLIP for the downstream task, CLIP applies the hand-crafted template "a photo of a {}” to extract the general class-level textual embedding, defined as $\mathbf{W}^{clip}=\{\mathbf{w}^{clip}_{i}\}_{i=1}^{{N}_{c}}$, where $\mathbf{w}^{clip}_{i}$ is the textual embedding of $i$-th class, and $N_{c}$ is the number of classes. Given the ‘class-name’ of $i$-th class, Word Embedded $e(\cdot)$ first embeds the hand-crafted description into a vectorized textual tokens: $\mathbf{t}^{clip}_{i}=e$("a photo of a {class-name}"). After that, Text Encoder $\psi$ maps the vectorized textual tokens $\mathbf{t}^{clip}_{i}$ to the class-level embedding: $\mathbf{w}^{clip}_{i}=\psi(\mathbf{t}^{clip}_{i})$.
@@ 3.2 Self-Enhanced Prompt Tuning
where $\theta_{l+1}$ is the (l+1)-th layer in Visual Encoder.
Importantly, the proposed SEP can also be easily applied to the Text Encoder for textual prompt tuning. When performing SEP on the Text Encoder, the representative tokens are selected by the front top-$L_{t}$ tokens containing the discriminative knowledge based on the hand-crafted templates "a photo of a {class-name}".
$\mathcal{L}_{kg\_v}=||\mathbf{\hat{f}}-\mathbf{f}||^{2}_{2}.$ (9)
By considering Eq. (3), the final objective is:
$\mathcal{L}=\mathcal{L}_{ce}(\mathbf{\hat{f}},\mathbf{W}^{sep})+\omega_{t}\mathcal{L}_{kg}(\mathbf{W}^{clip},\mathbf{W}^{sep})+\omega_{v}\mathcal{L}_{kg\_v}(\mathbf{\hat{f}},\mathbf{f})+\mathcal{L}_{ce\_v}(\mathbf{\hat{f}}),$ (10)
4 Experiments
4.1 Experimental Setup
Training Details. Our implementation is adapted from the public code [52; 45] based on the CLIP with the backbone of ViT-B/16 [8]. The prompt’s length $L_{v}$ and $L_{t}$ is set as 4 and 6 for the visual prompt tuning and textual prompt tuning. The hand-crafted templates initialized with "X X X X X X { }" are used for the textual prompt tuning. The final performance is averaged over three standard random seeds(1/2/3). Adam optimizer is applied for optimization with the learning rate of 2.5e-3 and the batch size of 32 up to the training epochs is 50. All experiments are conducted on RTX 3090.
@@ 4.2 Comparison with Existing methods
[…]
ImageNet -V2 -S -A -R Avg.
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
ProGrad 72.24 64.73 47.61 49.39 74.58 59.08
KgCoOp 71.2 64.1 48.97 50.69 76.7 60.12
MaPLe 70.72 64.07 49.15 50.9 76.98 60.27
DAPT 71.67 64.5 49.53 51.1 76.33 60.37
SEP 71.03 64.30 49.75 50.70 77.43 60.54
Table 3: Comparison of domain generalization from ImageNet to its variants.
[…]
Table 4: Effect of $\omega_{v}$.
4.3 Ablation Study
@@
[…]
@@
Table 6: Effect of Self-enhanced Prompt tuning.
[…]
@@
Table 7: Effect of Multi-modal Prompt tuning.
@@
TFM 85.96 76.53 80.97
Table 10: Effect of inserting SEP into different layers.
Layers {4} {8} {10} {4,8} {4,8,10} {3,6,9} {2,4,6,8,10} {1,2,3,4,5,6,7,8,9,10,11}
Base 84.66 84.62 85.16 85.38 85.44 85.32 86.27 85.98
New 74.8 76.07 75.34 75.19 75.41 74.95 75.92 76.49
H 79.43 80.12 79.95 79.96 80.11 79.80 80.76 80.96
Effect of Template Initialization In the proposed SEP, we initialized the templates with ‘X X X X {}’. We thus make a comparison with the hand-crafted templates, such as ‘a photo of {}’, ‘this is a photo {}’, and summarize the related results in Table 11. It can be observed that the proposed methods is less sensitive to the initialization of the template for the textual prompt tuning. The reason is that the self-enhanced prompt is major affected by the pre-trained tokens, having less relation to the template.
Table 11: Effect of Template Initialization
Templates Base New H
‘a photo of a{}’ 85.58 75.89 80.44
‘this is a photo {}’ 85.53 75.92 80.44
‘X X X X {}’ 85.56 75.90 80.44
Effect of Prompt length In this section, we analyze the effect of prompt length for the Self-Enhanced Prompt tuning, e.g., $L_{v}$ and $L_{t}$ for the visual and textual prompt tuning, respectively. As shown in Table 12, we can observe that the proposed SEP is less sensitive to the prompt length. For example, for setting SEP with the same $L_{t}$ for textual prompt tuning, different methods obtain similar performance, which is also the same for the prompt length for the visual prompt.
Table 12: Effect of Prompt Length for $L_{t}$ and $L_{v}$.
$L_{t}$ 2 2 2 2 4 4 4 4 6 6 6 6 8 8 8 8 16 16 16 16
$L_{v}$ 2 4 6 8 2 4 6 8 2 4 6 8 2 4 6 8 2 4 6 8
Base 85.38 85.43 85.45 85.56 85.56 85.56 85.62 85.51 85.87 85.98 85.88 85.85 85.91 85.92 85.93 85.88 85.99 85.94 85.94 85.91
New 76.27 75.91 75.99 75.82 75.82 75.90 75.82 75.94 76.22 76.49 75.98 75.92 75.48 75.92 75.89 75.84 76.35 76.02 76.03 76.24
H 80.57 80.38 80.44 80.40 80.40 80.44 80.42 80.44 80.76 80.96 80.63 80.58 80.36 80.61 80.60 80.55 80.88 80.68 80.68 80.79

### v3

# SEP: Self-Enhanced Prompt Tuning for Visual-Languge Model

@@ Abstract
Prompt tuning based on Context Optimization (CoOp) effectively adapts visual-language models (VLMs) to downstream tasks by inferring additional learnable prompt tokens. However, these tokens are less discriminative as they are independent of the pre-trained tokens and fail to capture input-specific knowledge, such as class-aware textual or instance-aware visual knowledge. Leveraging the discriminative and generalization capabilities inherent in pre-trained tokens, we introduce a novel approach named Self-Enhanced Prompt Tuning (SEP). The core principle of SEP involves adapting the learnable prompt tokens at each encoder layer from the corresponding self-pretrained tokens, thereby explicitly incorporating discriminative prior knowledge to enhance both textual-level and visual-level embeddings. Furthermore, SEP’s self-enhanced tokens not only boost discrimination but also mitigate domain shifts in unseen domains, enhancing generalization. In practice, SEP selects several representative tokens from all pre-trained tokens for each input data at every layer of the text/visual encoders. Subsequently, a Token Fusion Module (TFM) is introduced to generate a self-enhanced token by merging these representative tokens with the learnable tokens using a cross-attention mechanism. This self-enhanced token is then concatenated with all pre-trained tokens, serving as input for subsequent encoder layers to produce the relevant embeddings. Comprehensive evaluations across various benchmarks and tasks confirm SEP’s efficacy in prompt tuning. Code: https://github.com/htyao89/SEP/.
@@ 1 Introduction
Recent studies demonstrate that the Visual Language Model (VLM) possesses a robust generalization capability in a variety of downstream tasks [2, 33]. However, the direct application of VLMs in scenarios with limited images is challenging because of their extensive parameters and the substantial image datasets required for training. Prompt tuning has recently emerged as an effective and straightforward method for leveraging the general knowledge of VLMs in specialized computer vision and machine learning tasks such as zero-shot learning, domain adaptation, domain generalization, and few-shot learning.
@@
Prompt tuning combines additional learnable tokens with pre-trained ones to generate a discriminative embedding, termed Context Optimization (CoOp) [52]. Recent advances in CoOp-based methods [52, 45, 4, 31, 17, 5, 22, 18, 41, 54] aim to develop input-relevant prompt tokens to enhance the discriminative power of embeddings, as illustrated in Figure 1(a). The term ‘input-irrelevant prompt’ refers to a prompt initialized random that does not depend on specific input-related knowledge. However, since these prompts are derived from the training domain, they may exhibit domain shifts from unseen domains, potentially degrading performance. In addition, image-conditional prompts, which incorporate image features into the learnable tokens, have been proposed in [51, 48] (see Figure 1(b)). Despite their specificity, the image-conditional prompts have a limited ability to enhance the class-level classifier. In summary, while existing CoOp-based methods generate learnable tokens that are independent of input knowledge, they struggle to utilize essential input-aware prior knowledge, such as instance-aware visual knowledge or class-aware textual knowledge. Therefore, how to incorporate discriminative and generalizable input-aware knowledge into learnable prompts is crucial to enhancing their effectiveness for prompt tuning.
@@ 2 Related Works
Vision-Language Models: Recently, Vision-Language Models (VLMs) [33, 2], which utilize both visual and textual encoders, have been trained on a large-scale dataset of image-text pairs, demonstrating significant generalization and discrimination capabilities. To further enhance the descriptive power of VLMs, improvements have been made in several areas: (1) employing more robust text or visual encoders [42, 49, 25]; (2) integrating visual and textual knowledge more deeply [23, 39]; and (3) increasing the volume of images used [33, 16, 36, 37]. To enhance the diversity of text descriptions, Masked Language Modeling (MLM) [20, 27] has been employed, which randomly omits words in the text to aid in representation learning. In contrast to MLM, masked autoencoder-based methods [13] have been introduced to improve descriptive capabilities by randomly masking image patches. Among the VLM frameworks, CLIP stands out as a straightforward yet effective model that uses contrastive loss to train separate visual and text encoders based on 400 million image-text pairs. Given its robust generalization, many CoOp-based methods have been developed using CLIP to adapt pre-trained VLMs for downstream tasks. In this work, we have implemented a prompt-tuning strategy to generate task-specific embeddings.
@@
Prompt Tuning: To adapt pre-trained Vision-Language Models (VLMs) to downstream tasks, various strategies involving prompt tuning have been employed [10, 31, 22, 48, 33, 24]. Initially, CLIP[33] uses a hand-crafted template ‘a photo of a [CLASS]’ for zero-shot prediction. However, these hand-crafted prompts often fail to adequately capture the nuances of downstream tasks, leading to the development of textual prompt tuning strategies that use learnable textual tokens alongside class tokens to enhance textual embeddings. For example, the hand-crafted prompts are replaced with the learnable prompt tokens in Context Optimization(CoOp) [52]. Moreover, Conditional Context Optimization(CoCoOp) [51] and VPT [48] generate an image-conditional prompt fusing the image feature and the learnable textual prompt to boost generalization in the unseen domain. To better use the essential knowledge contained in frozen CLIP, Knowlege-Guided Context Optimization(KgCoOp) [45], ProGrad [53], and Prompt Regularization(ProReg) [54] all construct a novel constraint term by constraining the consistency between the learnable knowledge and the essential general knowledge. Unlike the above methods, which consider textual prompts, Ensembling Context Optimization(ECO) [1] employs prompt ensembling to combine multiple prompts. Furthermore, ProDA [28] considers the prompt’s prior distribution learning, and Distribution-Aware Prompt Tuning (DAPT) [5] optimizes the learnable prompt by maximizing inter-dispersion. Knowlege-Aware Prompt Tuning(KAPT) [17] employs external knowledge to boost generalization on unseen classes. PLOT [4] applies optimal transport to match the vision and text modalities for generating the discriminative and visual-aligned local textual prompt tokens. Besides the textual prompt tuning, Multi-modal Prompt Learning (MaPLe) [18] and PromptSRC [19] conduct the visual-textual prompt tuning by jointly conducting the prompt tuning on the visual and text encoders. Multitask Vision-Language Prompt Tuning (MVLPT) [38] integrates cross-task knowledge, enhancing the versatility of prompt tuning in VLMs. Finally, DenseCLIP [34] and CLIP-Adapter [12] introduce context-aware strategies and adapters to refine the embeddings for specific tasks, further illustrating the dynamic evolution of prompt tuning in the field of vision-language models. Recently, GraphAdaper [26] is proposed to fuse visual and textual knowledge with two Graph-based Adapters, and the Class-aware prompt is proposed in TCP [46] to inject class-level knowledge into the prompt.
@@ 3.1 Preliminaries
The frozen Contrastive Language-Image Pre-training(CLIP) trained with a large number of images exhibits robust generalization capacity to new classes or images, e.g., CLIP shows impressive performance on zero-shot learning and domain generalization tasks, Therefore, existing prompt tuning based on Context Optimization(CoOp) is proposed based on CLIP. Given an image along with all class names, CLIP extracts the visual and text embeddings with the visual and text encoders, and the contrast loss is calculated to align those two embeddings. To effectively adapt CLIP for the downstream task, CLIP applies the hand-crafted template ”a photo of a {}” to extract the general class-level textual embedding, defined as $\mathbf{W}^{clip}=\{\mathbf{w}^{clip}_{i}\}_{i=1}^{{N}_{c}}$, where $\mathbf{w}^{clip}_{i}$ is the textual embedding of $i$-th class, and $N_{c}$ is the number of classes. Given the ‘class-name’ of $i$-th class, Word Embedded $e(\cdot)$ first embeds the hand-crafted description into a vectorized textual tokens: $\mathbf{t}^{clip}_{i}=e$(”a photo of a {class-name}”). After that, Text Encoder $\psi$ maps the vectorized textual tokens $\mathbf{t}^{clip}_{i}$ to the class-level embedding: $\mathbf{w}^{clip}_{i}=\psi(\mathbf{t}^{clip}_{i})$.
@@ 3.2 Self-Enhanced Prompt Tuning
where $\theta_{l+1}$ is the (l+1)-th layer in Visual Encoder.
Importantly, the proposed SEP can also be easily applied to the Text Encoder for textual prompt tuning. When performing SEP on the Text Encoder, the representative tokens are selected by the front top-$L_{t}$ tokens containing the discriminative knowledge based on the hand-crafted templates ”a photo of a {class-name}”.
$\mathcal{L}_{kg\_v}=||\mathbf{\hat{f}}-\mathbf{f}||^{2}_{2}.$ (9)
By considering Eq. (3), the final objective is:
$\mathcal{L}=\mathcal{L}_{ce}(\mathbf{\hat{f}},\mathbf{W}^{sep})+\omega_{t}\mathcal{L}_{kg}(\mathbf{W}^{clip},\mathbf{W}^{sep})$ (10)
$+\omega_{v}\mathcal{L}_{kg\_v}(\mathbf{\hat{f}},\mathbf{f})+\mathcal{L}_{ce\_v}(\mathbf{\hat{f}}),$ (11)
4 Experiments
Training Details. Our implementation is adapted from the public code [52, 45] based on the CLIP with the backbone of ViT-B/16 [8]. The prompt’s length $L_{v}$ and $L_{t}$ is set as 4 and 6 for the visual prompt tuning and textual prompt tuning. The hand-crafted templates initialized with ”X X X X X X { }” are used for the textual prompt tuning. The final performance is averaged over three standard random seeds(1/2/3). Adam optimizer is applied for optimization with the learning rate of 2.5e-3 and the batch size of 32 up to the training epochs is 50. All experiments are conducted on RTX 3090.
@@ 4.1 Comparison with Existing methods
Table 2: Comparison of domain generalization.
ImageNet -V2 -S -A -R Avg.
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
ProGrad 72.24 64.73 47.61 49.39 74.58 59.08
KgCoOp 71.2 64.1 48.97 50.69 76.7 60.12
MaPLe 70.72 64.07 49.15 50.9 76.98 60.27
DAPT 71.67 64.5 49.53 51.1 76.33 60.37
SEP 71.03 64.30 49.75 50.70 77.43 60.54
@@
4.2 Ablation Study
Table 5: Effect of Self-enhanced Prompt tuning.
@@ 4.2 Ablation Study
Table 6: Effect of inserting SEP into different layers.
Layers {4} {8} {10} {4,8} {4,8,10} {3,6,9} {2,4,6,8,10} {1,2,3,4,5,6,7,8,9,10,11}
Base 84.66 84.62 85.16 85.38 85.44 85.32 86.27 85.98
New 74.8 76.07 75.34 75.19 75.41 74.95 75.92 76.49
H 79.43 80.12 79.95 79.96 80.11 79.80 80.76 80.96
Table 7: Effect of Multi-modal Prompt tuning.
@@
TFM 85.98 76.49 80.96
Figure 3: Effect of $W_{v}$.